Q: Assistance in understanding and solving this example on Probability and Statistical with the steps of the solution to better understand, thanks.

**Please give the step by steps with details to completely see how the solution came about especially on the table.

1) Let X1 and X2 be two independent random variables. Each follows a discrete Uniform distribution on S={1,2,3,4}. Let X=X1 and let Y=X1+X2.

(a) Fill in the joint probability distribution f(x,y) in the table:

x\y   2    3    4    5    6    7    8

1-------------------------------------------

2-------------------------------------------

3-------------------------------------------

4 ------------------------------------------

(b) Find the marginal pmfs of X.

(c) Find the marginal pmf of Y.

(c ) Ca1culate Cov(X, Y).
Answer #1

for x is discrete uniform ; therefore P(X=x)=1/4

hence a) below is joint probability distribution f(x,y) in the table :

y
x 2 3 4 5 6 7 8 Total
1    1/16    1/16    1/16    1/16 0        0        0           1/4
2 0           1/16    1/16    1/16    1/16 0        0           1/4
3 0        0           1/16    1/16    1/16    1/16 0           1/4
4 0        0        0           1/16    1/16    1/16    1/16    1/4
Total    1/16    1/8    3/16    1/4    3/16    1/8    1/16 1.0000

b)

marginal pmf of X:

x P(x)
1    1/4
2    1/4
3    1/4
4    1/4

c)

marginal pmf of Y:

y P(y)
2    1/16
3    1/8
4    3/16
5    1/4
6    3/16
7    1/8
8    1/16

d)

E(X)=\sumxP(x)=1*1/4+2*1/4+3*1/4+4*1/4=2.5

E(Y)=\sumyP(y)=2*1/16+3*1/8+4*3/16+5*1/4+6*3/16+7*1/8+8*1/16 =5

E(XY)=\sumxyP(x,y)=1*2*1/16+1*3*1/16+...+4*8*1/16=13.75

therefore Cov(X,Y)=E(XY)-E(X)*E(Y)=1.25